OpenStudy (anonymous):

The equations 5x + 2y= 48 and 3x + 2y= 32 represent the money collected from school concert tickets sales during two class periods. If x represent the cost for each adult ticket and y represent the cost for each student ticket, what is the cost for each adult ticket ? A 20 B 10 C 8 D 4

OpenStudy (anonymous):

Need help

OpenStudy (anonymous):

The equations 5x+2y=48 and 3x+2y=32 represent the money collected from school concert ticket sales during two class periods. If x represents the cost for each adult ticket and y represents the cost for each student ticket, what is the cost for each adult ticket? . You basically have two unknowns and two equations (system of equations). You can use the "addition method" to solve. 5x+2y=48 3x+2y=32 . Multiply both sides of the second equation by -1 5x+2y = 48 -3x-2y =-32 . Now, add the two equations together: 5x+2y = 48 -3x-2y =-32 -------------- 2x = 16 dividing both sides by 2: x = $8 (adult ticket cost)

OpenStudy (anonymous):

use the subtraction method that way u don't need to multipy 5x+2y=48 -3x+2y=32 ____________ 2x=16 x=16/2=8

OpenStudy (anonymous):

Thank you